Please help - I don't understand.
5/x+6 + 2/x^2+7x+6 = 3/x+1
In long form it is... 5 over x plus 6, plus 2 over x squared plus 7x plus 6, equals 3 over x plus 1.
Thank you so much!!

Answer by stanbon(75887)

5/x+6 + 2/x^2+7x+6 = 3/x+1
Multiply thru by x^2+7x+6=(x+6)(x+1) to get:
5(x+1)+2 = 3(x+6)
5x+5+2=3x+18
2x=11
x=11/2
Cheers,
Stan H.